Q10 1 Point Assume that the following sequence converges. Find its limit. 2, 2/2, ( 2/2/2, ... N | ооооо
Answer #1

35, 3/232, 3/23/2372 Let the final value of above series is Let = 3/23/23/1... x=3/2xx x² = 21 =) L=0 & x²= 2 x = 2 2 possibl